Direct Labor Variances

The following data relate to labor cost for production of 4,600 cellular telephones:

Actual: 3,080 hrs. at $15.4

Standard: 3,030 hrs. at $15.7

a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Rate variance $

Time variance $

Total direct labor cost variance $
Answer #1

Rate variance

= ACTUAL HRS ( ACTUAL RATE - STANDARD RATES )

= 3080 ( 15.4 - 15.7)

= $924 F.

Time variance

= Std. Rate ( actual hrs - std. Hrs)

= 15.7 ( 3080 - 3030 )

$785 adv.

C. Total direct labour cost variance

= SR × SH - Ar × AH

= 15.7 × 3030 .- 15.4 × 3080

=$47571 - $ 47432

= $139 fAvourable